# Sparse Feature Coactivation Reveals Causal Semantic Modules in Large Language Models

### v2

@@ 3.3 Component Organization
Having established the causal role and composability of concept and relation components, we next analyze their distribution across network layers and the relative importance of nodes within these components. To quantify the causal importance of an individual concept node, we compute the average post-ablation KL divergence from the original output distribution across all relations. For a relation node, we compute the same average across all concepts. Figure 4 displays results for the country facts task, with more examples in Appendix J. Concept and relation components show distinct distribution patterns across network layers. 8 of 10 country components begin in the first layer, as do all word and verb components in other tasks. Some concept components span nearly all layers (e.g., for China and love), while others concentrate in early to middle layers (e.g., for Nigeria and red). In contrast, relation components concentrate in later layers, with most spanning only the last quarter to half of the network’s layers (e.g., for country facts and target translation languages), and a few spanning across the network (e.g., for verb transformations). This suggests that representations of more concrete entities are established earlier in processing, while more abstract relations emerge later. While concept component nodes exhibit various trends in KL divergence across layers, we find that relation components for the country facts task have stronger causal impact in later layers, while the opposite trend holds for word translation.
@@ 3.4 Single Feature Steering Comparison
We hypothesize that our approach of steering based on a group of features is advantageous over steering based on an individual feature (as done in some prior works, e.g., Anthropic, 2023; Templeton et al., 2024). To validate this, we compared its efficacy against a simpler baseline: intervening on only the single most causally impactful feature. Specifically, we identified this top feature for each concept and relation as the one whose individual ablation produced the highest KL divergence, then in line with earlier results, used it for ablation and steering. For the country facts task, the baseline respectively achieved average success rates of 83%, 83%, and 75% for composite country-relation steering. These results fall notably short of the success rates achieved using full components (respectively 96%, 93%, and 92%). This suggests that concepts are represented distributively across populations of features, and our method’s success comes from identifying and manipulating these entire functional groups. Neuroscience research has similarly found evidence for populations of firing neurons encoding information in human brains Xie et al. (2024).
3.5 Cross-Task Evaluation
Lastly, we explore whether ablating our identified components impacts performance on other tasks, an important question for the practicality of manipulating models based on them. Table 5 presents the model’s average accuracy for each task when our identified components are ablated. When ablating components from the country facts task, 100% accuracy (or near it) is maintained in most cases, even when ablating components for some country-fact pairs and prompting the model for other ones. Accuracy on translation degrades slightly, possibly due to the intersection of some country- and language-specific information. In ablating word translation components, accuracy remains high on other tasks, but degrades slightly on translation itself. This may be expected, given that the relations in this task are of highly similar nature, and even overlap by a small number of features. In ablating verb transformation components, we see moderately degraded accuracy for all tasks to as low as 64-73%. This suggests that the components for this task are less specific than other tasks, in line with our relatively low steering performance with them. Nonetheless, this demonstrates that extracted components are generally fairly specific to the functions we associate them with, promising for manipulating deployed models with them.
Ablated Task C.F. Acc. W.T. Acc. V.T. Acc.
@@ 3.5 Cross-Task Evaluation
Word Translation 1.00 0.83 0.99
Verb Transformation 0.73 0.64 0.66
@@ 4 Related Work
Prior work has applied various methods to interpret large neural networks as compositions of subnetworks that perform specific functions (Olah et al., 2020; Elhage et al., 2021; Sharkey et al., 2025). Specialized circuits have been discovered for various functions of LLMs, including in-context learning (Olsson et al., 2022), numerical comparison (Hanna et al., 2023), indirect object recognition (Wang et al., 2023), and broader natural language fluency AlKhamissi et al. (2025). More recent work has developed automated circuit discovery methods to improve scalability (Goldowsky-Dill et al., 2023; Conmy et al., 2023; Hsu et al., 2025; Bhaskar et al., 2024). However, these approaches can be computationally expensive and yield circuits that are difficult to interpret.
@@
As large neural networks like LLMs are thought to encode sparse features across polysemantic neurons (Elhage et al., 2022), dictionary learning techniques like SAEs promise greater interpretability by extracting monosemantic features from them (Bricken et al., 2023; Huben et al., 2024). This has allowed automated circuit discovery approaches to operate on these features instead of neurons (Marks et al., 2025), but high computational costs remain a barrier. Like our work, Li et al. (2025) used SAE feature coactivation patterns to analyze their geometry, finding spatial clustering of related concepts. Transcoders (Dunefsky et al., 2024) extend SAEs by learning to approximate layerwise computations. Cross-layer transcoders comprise the Anthropic circuit tracing tool (Ameisen et al., 2025), which visualizes LLM responses with sparse but still difficult to interpret computation graphs. Building on these efforts, we instead follow Li et al. to leverage SAE feature coactivation, applying node pruning based on activation density to automatically discover semantically coherent, context-consistent components that can influence LLMs’ relational reasoning outputs individually and in combination. Our approach offers an efficient framework for analyzing and controlling LLM behavior without exhaustive circuit tracing.
@@
Orthogonally, factual knowledge is believed to reside in the feedforward layers of transformer-based LLMs. Geva et al. (2021) characterized these layers as key-value memories mapping textual patterns to vocabulary distributions. Dai et al. (2022) identified “knowledge neurons” whose activations correlate with specific facts. These insights informed approaches for editing factual knowledge stored in LLMs (De Cao et al., 2021; Mitchell et al., 2022; Meng et al., 2022). Geva et al. (2023) described the recall of factual associations as a three-step process of subject enrichment, relation propagation, and attribute extraction, while Hernandez et al. (2024) demonstrated that relation decoding in transformers can be approximated by simple linear transformations. Recent work on “knowledge circuits” has begun tracing causal pathways underlying factual recall (Yao et al., 2024; Ou et al., 2025). Most relevant to this work, Merullo et al. (2024) demonstrated that LLMs implement word2vec-style vector arithmetic through their feedforward layers to solve some relational tasks. Our work extends this effort by identifying finer-grained modular groups of features that represent human-interpretable concepts and relations which are similarly composed to form responses.
@@ 5 Conclusion
By tracing sparse feature coactivation patterns across layers, we uncover semantically coherent, context-consistent components that encode both concrete entities and concepts as well as abstract relations. We find that ablating or amplifying these components reliably alters model behavior. Moreover, composing country and relation components induces compound counterfactual responses. The layer-wise distribution of components reveals a hierarchical organization: entity features emerge in early layers, while relational features cluster in deeper layers. Overall, our method provides a lightweight, interpretable framework for analyzing and steering LLM behavior without full-scale circuit tracing. This work advances scalable interpretation and control of LLM behavior by leveraging pretrained SAEs to efficiently identify functional components. Given the increased adoption of LLMs by humans on a greater variety of tasks, such capabilities are vital for sustaining a human-machine enterprise that is not only reliable and safe, but also transparent, aligned with human goals, and resilient to misuse and misinterpretation.
@@
Some key limitations of our study warrant acknowledgment:
@@
First, due to computational constraints, our investigation was constrained to only 3 multi-relational reasoning tasks with small numbers of examples. When it comes to the task space, we attempted to add some additional tasks around Spanish verb conjugation, factual information about Nobel Prize winners, and taxonomical information about animals, [^10] More discussion about animal taxonomy results in Appendix K. but Gemma 2 2B could not perform these tasks accurately off the shelf (for the former, this is likely due to Gemma not being trained on much Spanish data), thus we would not expect to identify semantic modules for these tasks. That said, the tasks we explored in this paper are diverse, representing factual and lexicosemantic knowledge, and similar forms of these tasks are also considered in prior relevant works Merullo et al. (2024); Todd et al. (2024); Hernandez et al. (2024). As such, we expect that our analytical framework generalizes to other domains.
@@
This is also nontrivial due to the requirement of manually selecting components based on various heuristics outlined in Section 3. While we believe it may be possible to automate these heuristics or even optimize component selection in future work, this would still take a significant amount of time and computational power. Nevertheless, we believe the examples in our datasets comprise sufficiently representative samples of the domains, e.g., countries around the world, animals from various classes, orders, and families, words of various syntactic categories, and verbs with both common and irregular past tense transformations.
@@
Computational constraints also restricted our analysis to standard JumpReLU sparse autoencoders (Rajamanoharan et al., 2024b), which can suffer from inconsistent feature quality, poor reconstruction, and weak functional alignment. These limitations have motivated various improved architectures (Braun et al., 2024; Rajamanoharan et al., 2024a, b; Dunefsky et al., 2024), which may yield even more refined functional module identification under our approach. As a proof-of-concept, we extended our translation analysis to Gemma 2 2B transcoders (Dunefsky et al., 2024), successfully rediscovering some causal features for translation languages that Hanna et al. manually select, and achieving about 27% steering success rate with our extracted components. [^11] See Appendix L for more details. This provides promise that this approach can work with broader interpretability methods than SAEs.
@@ Appendix B Full Task Details
In this appendix, we provide the full list of concepts included in our analysis for each task. Additionally, as our accuracy and steering success evaluations importantly accounted for multiple possible correct answers to prompts, we provide lists of correct answers for each explored task’s concept-relation pairs below. These lists were refined through inspecting correct answers counted as incorrect in intermediate results. Importantly, the off-the-shelf tested models achieve 100% accuracy on all tasks under the few-shot prompts used to identify components, as well as the zero-shot prompts used for ablation and steering experiments. The model’s inconsistent performance may be another reason for the relatively low steering success rates on this task.
@@ B.1 Country Facts
    • China $\rightarrow$ Beijing
@@
    • Russia $\rightarrow$ Ruble, , RUB
@@ B.2 Word Translation
  • Parts of speech, e.g., poisson (fish) and pêcher (to fish) as correct French translations of fish)
@@
    • cat $\rightarrow$ gato, gata, felino
    • dog $\rightarrow$ perro, perra
@@
    • here $\rightarrow$ aquí, acá, aqui, Aqui
@@
    • love $\rightarrow$ amar, amo, amas, aman, encanta
@@
    • run $\rightarrow$ correr, corro, corres, corren, corrida, corre
    • you $\rightarrow$ tú, eres
@@
    • beautiful $\rightarrow$ belle, beau, bel
    • cat $\rightarrow$ chat, chatte, félin
    • dog $\rightarrow$ chien, chienne
@@
    • run $\rightarrow$ courir, cours, course, couler, coureur, A pied
@@
    • beautiful $\rightarrow$ schön
    • cat $\rightarrow$ Katze
    • dog $\rightarrow$ Hund
@@
    • good $\rightarrow$ gut, gute, guten, Gut
@@ B.3 Verb Transformation
    • like $\rightarrow$ dislike, hate, avoid
@@
    • sink $\rightarrow$ float, rise, ascend
@@ Appendix G Steering Details and Hyperparameters
  • Country facts: $\alpha_{c}=0.1$, $\alpha_{r}=0.45$, $\alpha^{\prime}_{c}=0.15$, and $\alpha^{\prime}_{r}=0.45$
  • Word translation: $\alpha_{c}=0.05$, $\alpha_{r}=0.45$, $\alpha^{\prime}_{c}=0.05$, and $\alpha^{\prime}_{r}=0.4$
  • Verb transformation: $\alpha_{c}=0.05$, $\alpha_{r}=0.45$, $\alpha^{\prime}_{c}=0.055$, and $\alpha^{\prime}_{r}=0.5$

### v3

@@ 3.3 Component Organization
Having established the causal role and composability of concept and relation components, we next analyze their distribution across network layers and the relative importance of nodes within these components. To quantify the causal importance of an individual node in a concept component, we compute the average post-ablation KL divergence from the original output distribution across all relations. For nodes in relation components, we compute the same average across all concepts. Figure 4 displays results for the country facts task, with more examples in Appendix K.
Concept and relation components show distinct distribution patterns across network layers. 8 of 10 country components begin in the first layer, as do all word/verb components in other tasks. Some concept components span nearly all layers (e.g., for China and love), while others concentrate in early to middle layers (e.g., for Nigeria and red). In contrast, relation components concentrate in later layers, with most spanning only the last quarter to half of the network’s layers (e.g., for country facts and target translation languages), and some spanning further across the network while having greater densities of nodes in later layers (e.g., for verb transformations). This may suggest that representations of more concrete entities are established earlier in processing, while more abstract relations emerge later. Concept component nodes exhibit various trends in KL divergence across layers. Meanwhile, relation components for country facts, past tense, and first letter have stronger causal impact in later layers, and the opposite trend holds for translation languages and capitalization.
@@ 3.4 Single Feature Steering Comparison
We hypothesize that our approach of steering based on a group of features is advantageous over steering based on an individual feature (as done in some prior works, e.g., Anthropic, 2023; Templeton et al., 2024). To validate this, we compared its efficacy against a simpler baseline: intervening on only the single most causally impactful feature. Specifically, we identified this top feature for each concept and relation as the one whose individual ablation produced the highest KL divergence, then in line with earlier results, used it for ablation and steering.
For the country facts task, the baseline respectively achieved average success rates of 83%, 83%, and 75% for concept, relation, and composite steering. These results fall notably short of the success rates achieved using full components (respectively 96%, 93%, and 90%). This suggests that concepts are represented distributively across populations of features, and our method’s success is largely attributable to identifying and manipulating these entire functional groups.
3.5 Specificity Evaluation
Lastly, we explore the degree to which ablating our identified components impacts performance on other tasks, an important question for the specificity of mechanisms and practicality of manipulating models based on them. Table 5 presents the model’s average accuracy for each task when our identified components are ablated. When ablating components from the country facts task, 100% accuracy (or near it) is maintained in most cases, even when ablating components for some country-fact pairs and prompting the model for other ones. Accuracy on translation degrades slightly, possibly due to the intersection of some country- and language-specific information. In ablating word translation components, accuracy remains high on other tasks, but degrades slightly on translation itself. This may be expected, given that the relations in this task are of highly similar nature, and even overlap by a small number of features. In ablating verb transformation components, we see moderately degraded accuracy for all tasks to as low as 62-73%. This suggests that the components for this task are less specific than other tasks, in line with our relatively low steering performance with them. Nonetheless, this demonstrates that extracted components are generally fairly specific to the functions we associate them with, promising for manipulating deployed models with them.
Ablated Task Country Facts Acc. Word Transl. Acc. Verb Transf. Acc.
@@ 3.5 Specificity Evaluation
Word Transl. 1.00 0.83 1.00
Verb Transf. 0.73 0.64 0.63
@@ 4 Related Work
Circuits and sparse feature dictionaries.
@@
Prior work has applied various methods to interpret large neural networks as compositions of subnetworks that perform specific functions (Olah et al., 2020; Elhage et al., 2021; Sharkey et al., 2025). Specialized circuits have been discovered for various functions of LLMs, including in-context learning (Olsson et al., 2022), numerical comparison (Hanna et al., 2023), indirect object recognition (Wang et al., 2023), and broader natural language fluency AlKhamissi et al. (2025). More recent work has developed automated circuit discovery methods to improve scalability (Goldowsky-Dill et al., 2023; Conmy et al., 2023; Hsu et al., 2025; Bhaskar et al., 2024). However, these approaches can be computationally expensive and yield circuits that are difficult to interpret. As large neural networks like LLMs are thought to encode sparse features across polysemantic neurons (Elhage et al., 2022), dictionary learning techniques like SAEs promise greater interpretability by extracting monosemantic feature directions from them (Bricken et al., 2023; Huben et al., 2024). This has allowed automated circuit discovery approaches to operate on these features instead of neurons (Marks et al., 2025), but high computational costs remain a barrier. Like our work, Li et al. (2025) used SAE feature coactivation patterns to analyze their geometry, finding spatial clustering of related concepts. Standard SAEs can also suffer from inconsistent feature quality, poor reconstruction, and weak functional alignment, prompting various architectural and training improvements (Braun et al., 2024; Rajamanoharan et al., 2024a, b). Transcoders (Dunefsky et al., 2024) extend SAEs by learning to approximate layerwise computations. Cross-layer transcoders comprise the Anthropic circuit tracing tool (Ameisen et al., 2025), which visualizes LLM responses with sparse but still difficult-to-interpret computation graphs. Building on these efforts, we follow Li et al. to leverage SAE feature coactivation, applying node pruning based on activation density to automatically discover semantically coherent, context-consistent components that can influence LLMs’ relational reasoning outputs individually and in combination. Our approach offers an efficient framework for analyzing and controlling LLM behavior without exhaustive circuit tracing.
Knowledge organization in LLMs.
@@
Orthogonally, prior work has used various methods to study the organization of lexical and factual knowledge in LLMs. Factual knowledge is believed to reside in the feedforward layers of transformer-based LLMs. Geva et al. (2021) characterized these layers as key-value memories mapping textual patterns to vocabulary distributions. Dai et al. (2022) identified “knowledge neurons” whose activations correlate with specific facts. These insights informed approaches for editing factual knowledge stored in LLMs (De Cao et al., 2021; Mitchell et al., 2022; Meng et al., 2022). Geva et al. (2023) described the recall of factual associations as a three-step process of subject enrichment, relation propagation, and attribute extraction, while Hernandez et al. (2024) demonstrated that relation decoding in transformers can be approximated by simple linear transformations. Recent work on “knowledge circuits” has begun tracing causal pathways underlying factual recall (Yao et al., 2024; Ou et al., 2025).
Highly relevant to this work, recent work has explored the composition of knowledge in LLMs. Merullo et al. (2024) demonstrated that LLMs implement word2vec-style Mikolov et al. (2013) vector arithmetic through their feedforward layers to solve some relational tasks. Bayat et al. (2025) leveraged active SAE features in contrastive paired prompts to steer LLM behaviors, showing evidence of composability of behaviors. In line with our word translation results, Dumas et al. (2025) used activation patching to reveal language-agnostic concept representations in LLMs, such that the translation language and concept can be steered independently. Our work extends these efforts by identifying finer-grained modular groups of feature directions that represent human-interpretable concepts and relations and similarly compose to form responses, possibly from a single relational prompt.
@@ 5 Conclusion
By tracing sparse feature coactivation patterns across layers, we uncover semantically coherent, context-consistent components encoding both concrete entities and abstract relations. We find that ablating or amplifying these components reliably alters model behavior. Moreover, composing concept and relation components induces compound counterfactual responses. The layer-wise distribution of components reveals a hierarchical organization: entity features emerge in early layers, while relational features cluster in deeper layers. Overall, our method provides a lightweight, interpretable framework for analyzing and steering LLM behavior without full-scale circuit tracing. This work advances scalable interpretation and control of LLM behavior by leveraging pretrained SAEs to efficiently identify functional components. Given the adoption of LLMs by humans on a greater variety of tasks, such capabilities are vital for sustaining a human-machine enterprise that is not only reliable and safe, but also transparent, aligned with human goals, and resilient to misuse and misinterpretation.
@@
We acknowledge that our study is limited in the following ways:
@@
First, due to computational constraints, our investigation was constrained to only 3 multi-relational reasoning tasks with small numbers of examples. When it comes to the task space, we attempted to add some additional tasks around Spanish verb conjugation, factual information about Nobel Prize winners, and taxonomical information about animals, [^12] More discussion about animal taxonomy results in Appendix L. but Gemma 2 2B could not perform these tasks accurately off the shelf (for the former, this is likely due to Gemma not being trained on much Spanish data), thus we would not expect to identify semantic modules for these tasks. That said, the tasks we explored in this paper are diverse, representing factual and lexicosemantic knowledge, and similar forms of these tasks are also considered in prior relevant works Merullo et al. (2024); Todd et al. (2024); Hernandez et al. (2024). As such, we expect that our analytical framework generalizes to other domains.
@@
This is also nontrivial due to the requirement of manually selecting components based on various heuristics outlined in Section 3. While we believe it may be possible to automate component selection in future work (as described below), this would still take a significant amount of time and computational power. Nevertheless, we believe the examples in our datasets comprise sufficiently representative samples of the domains, e.g., countries around the world, animals from various classes, orders, and families, words of various syntactic categories, and verbs with both common and irregular past tense transformations.
@@
Computational constraints also restricted our analysis to standard JumpReLU sparse autoencoders (Rajamanoharan et al., 2024b), which can suffer from inconsistent feature quality, poor reconstruction, and weak functional alignment. These limitations have motivated various improved architectures (Braun et al., 2024; Rajamanoharan et al., 2024a, b; Dunefsky et al., 2024), which may yield even more refined functional module identification under our approach. As a proof-of-concept, we extended our translation analysis to Gemma 2 2B transcoders (Dunefsky et al., 2024), successfully rediscovering some causal features for translation languages that Hanna et al. manually select, and achieving about 27% steering success rate with our extracted components. [^13] See Appendix M for more details. This provides evidence that this approach can work with broader interpretability methods than SAEs.
@@
Reliance on manual intervention.
While our method of extracting components based on sparse feature coactivation is fully automated, as detailed in Section 3.1, we manually selected and/or combined components to represent each concept and relation using heuristics dependent on the task. This may give an impression that significant manual intervention is required, compromising the utility of our method. However, the focus of our work was primarily to automate the former step of component extraction, and given that, we found it relatively simple to identify semantically coherent and impactful components (the latter step). As we expanded the work to more tasks, we found that different choices maximized performance depending on the nature of the task, resulting in minor variations to our method.
That said, after conducting this work, we believe these heuristics can be replaced with a simple algorithm in future work. First, as our approach already does, we can apply our sparse feature coactivation method to identify several components for each concept-relation prompt, and sort them by their KL divergence of the LLM’s next token probability before and after ablation. Second, we can represent each unique concept and relation by identifying the most impactful component(s) which overlap among all prompts for that concept or relation, then taking the intersection of them. Another advantage of such an approach is that it removes dependence on feature descriptions, e.g., those we used from Neuronpedia.
@@
Acknowledgments
We thank our anonymous reviewers for their thoughtful and constructive feedback. This research was supported in part through computational resources and services provided by Advanced Research Computing at the University of Michigan, Ann Arbor. Google Gemini 3 (gemini.google.com) and Anthropic Claude (claude.ai) were used for minor language edits and generation of boilerplate and visualization code.
@@ Appendix B Full Task Details
In this appendix, we provide the full list of concepts included in our analysis for each task. Additionally, as our accuracy and steering success evaluations importantly accounted for multiple possible correct answers to prompts, we provide lists of correct answers for each explored task’s concept-relation pairs below. During evaluation, LLMs generate an answer of no more than 5 tokens. An LLM’s answer is generally counted as correct if it contains any of the correct answers in the lists below, except for the following caveats.
In the first letter transformation and translation cases where correct answers are less than 4 characters long, the LLM must generate a correct answer on its own surrounded by whitespace or non-letters. This constraint reduces the possibility of mistakenly marking an LLM output that is not a correct answer but happens to contain one as correct, e.g., the string "te" appears in many Spanish and French words, but not all such words should be taken as correct answers for translating you. However, it is also common in translation for an LLM to output a token pertaining to a correct answer, then follow it up with tokens that morph its meaning (e.g., "Rotwein" when steering the LLM to output the German word for red). To account for this, in translation, we always mark LLM outputs beginning with a correct answer as correct. Evaluation is case-insensitive, except for the capitalization relation of the verb transformations task.
The below lists of correct answers were refined through inspecting correct answers counted as incorrect in intermediate results. Importantly, the off-the-shelf tested models achieve 100% accuracy on all tasks under the few-shot prompts used to identify components, as well as the zero-shot prompts used for ablation and steering experiments.
@@ B.1 Country Facts
    • China $\rightarrow$ Beijing, Peking
@@
    • Russia $\rightarrow$ Ruble, , RUB, руб
@@ B.2 Word Translation
  • Parts of speech, e.g., poisson (fish) and pêcher (to fish) as correct French translations of fish
@@
    • cat $\rightarrow$ gato, gata, gatos, gatas, felino
    • dog $\rightarrow$ perro, perra, perros, perras
@@
    • here $\rightarrow$ aquí, acá, aqui
@@
    • love $\rightarrow$ amar, amor, amo, amas, aman, encanta
@@
    • run $\rightarrow$ correr, corro, corres, corren, corrida, corre, carrera, corredor
    • you $\rightarrow$ tú, eres, usted, le, te, ustedes, vosotros
@@
    • beautiful $\rightarrow$ belle, beau, bel, beauté
    • cat $\rightarrow$ chat, chatte, chats, chattes, félin
    • dog $\rightarrow$ chien, chienne, chiens, chiennes
@@
    • run $\rightarrow$ courir, cours, course, couler, coureur
@@
    • beautiful $\rightarrow$ schön, schöne, schönen
    • cat $\rightarrow$ Katze, Katzen
    • dog $\rightarrow$ Hund, Hunde
@@
    • good $\rightarrow$ gut, gute, guten
@@ B.3 Verb Transformation
    • like $\rightarrow$ dislike, hate, avoid, unlike
@@
    • sink $\rightarrow$ float, rise, ascend, stand
@@ Appendix G Steering Details and Hyperparameters
  • Country facts: $\alpha_{c}=0.1$, $\alpha_{r}=0.4$, $\alpha^{\prime}_{c}=0.15$, and $\alpha^{\prime}_{r}=0.45$
  • Word translation: $\alpha_{c}=0.05$, $\alpha_{r}=0.4$, $\alpha^{\prime}_{c}=0.05$, and $\alpha^{\prime}_{r}=0.35$
  • Verb transformation: $\alpha_{c}=0.05$, $\alpha_{r}=0.65$, $\alpha^{\prime}_{c}=0.05$, and $\alpha^{\prime}_{r}=0.7$
For the verb transformation task, we additionally tried tuning a separate steering strength for each (target) relation. For individual concept steering, this made no difference in performance. For relation steering, past tense achieved up to 3.1%, capitalization 50.0%, and first letter 87.5%. However, since antonym and synonym remained at 0% steering success rate, choosing a single $\alpha_{r}$ in facilitating composite steering was not possible. As such, we chose to omit these results to avoid unnecessarily complicating our experimental paradigm. However, for future use of this method, it may be advantageous to explore more fine-grained selection of steering strengths for heterogeneous tasks like this one.
To provide more information about how performance varies with respect to steering strength, Tables 9, 10, and 11 list the individual concept/relation and composite steering accuracies across all searched steering strengths and tasks.
Steering Strength Country Facts Word Translation Verb Transformation
0.05 0.87 0.75 0.48
0.10 0.96 0.51 0.30
0.15 0.84 0.18 0.11
0.20 0.54 0.09 0.06
0.25 0.22 0.03 0.06
0.30 0.17 0.00 0.10
0.35 0.15 0.00 0.06
0.40 0.15 0.00 0.08
0.45 0.16 0.00 0.08
0.50 0.16 0.00 0.08
0.55 0.16 0.00 0.08
0.60 0.15 0.00 0.07
0.65 0.15 0.00 0.08
0.70 0.16 0.00 0.08
0.75 0.17 0.00 0.10
0.80 0.20 0.00 0.09
0.85 0.20 0.03 0.06
0.90 0.20 0.03 0.06
0.95 0.19 0.03 0.06
1.00 0.18 0.03 0.06
Table 9: Overall task accuracy for individual concept steering across candidate steering strengths. Accuracy under selected steering strength for each task in bold.
Steering Strength Country Facts Word Translation Verb Transformation
0.05 0.17 0.03 0.10
0.10 0.60 0.35 0.09
0.15 0.68 0.73 0.11
0.20 0.67 0.94 0.11
0.25 0.70 0.98 0.11
0.30 0.77 0.98 0.10
0.35 0.85 0.98 0.14
0.40 0.93 0.98 0.16
0.45 0.92 0.91 0.19
0.50 0.92 0.82 0.19
0.55 0.88 0.65 0.21
0.60 0.83 0.48 0.22
0.65 0.78 0.35 0.22
0.70 0.78 0.20 0.21
0.75 0.77 0.17 0.19
0.80 0.75 0.14 0.19
0.85 0.73 0.09 0.19
0.90 0.70 0.09 0.17
0.95 0.67 0.06 0.17
1.00 0.63 0.02 0.17
Table 10: Overall task accuracy for individual relation steering across candidate steering strengths. Accuracy under selected steering strength for each task in bold.
Steering Strengths Country Facts Word Translation Verb Transformation
0.00, 0.35 0.01
0.00, 0.40 0.01
0.00, 0.45 0.01
0.00, 0.60 0.15
0.00, 0.65 0.16
0.00, 0.70 0.16
0.05, 0.35 0.74 0.64
0.05, 0.40 0.78 0.59
0.05, 0.45 0.77 0.48
0.05, 0.60 0.17
0.05, 0.65 0.18
0.05, 0.70 0.19
0.10, 0.35 0.85 0.53
0.10, 0.40 0.85 0.47
0.10, 0.45 0.86 0.36
0.10, 0.60 0.16
0.10, 0.65 0.18
0.10, 0.70 0.18
0.15, 0.35 0.86
0.15, 0.40 0.88
0.15, 0.45 0.90
Table 11: Overall task accuracy for composite concept-relation steering across candidate steering strength pairs. Accuracy under selected steering strength pair for each task in bold.